AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 243     Entry time: Tue Feb 22 20:08:40 2022
Author: David, Yuri 
Subject: Current status 

Recently the pressure started to raise a bit, we have increased the temperature from 129.1 K to 129.2 K. Current pressure at S1 2.10e-7.

Merger, Tape server (WS3) checks: DONE, writing ~1 MB/s, refreshing is fine

Leakage current (WS4): 3.46 uA , stable current

DAQ (WS2): temperature checked, clock status:ok, adc calibration:ok, white rabbit check:ok

DAQ (WS2): everything ok

Online monitoring (WS 5): done

Attachment 1: 2022-02-22_21-14-23exrate.png  156 kB  | Hide | Hide all
2022-02-22_21-14-23exrate.png
Attachment 2: 2022-02-22_21-14-01xytotal.png  197 kB  | Hide | Hide all
2022-02-22_21-14-01xytotal.png
Attachment 3: 2022-02-22_21-13-11extotal.png  165 kB  | Hide | Hide all
2022-02-22_21-13-11extotal.png
ELOG V3.1.4-unknown